For the curve y = x^2 + x, at what point does the normal line at (0,0) intersect the tangent line at (1,2)?
Answer by Alan3354(69443)

For the curve y = x^2 + x, at what point does the normal line at (0,0) intersect the tangent line at (1,2)?
-------------
The tangent line at (0,0) is y = x --> normal line is y = -x
---
The tangent line at (1,2) is y = 3x-1
---
y = -x
y = 3x - 1
-x = 3x - 1
x = 1/4
etc